Prev Next

SW&IT Meeting Minutes 2022-01-31

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220124
LR20220124 - Bec, Matthieu 40.0
LR20220124 - Mechtley, Brandon 40.0
LR20220124 - Eychaner, Glenn 0.0
LR20220124 - Edmeades, David 40.0
LR20220124 - Grenz, Paul 20.0
LR20220124 - Hooper, Stephen 0.0
LR20220124 - Zhang, Yang 40.0
LR20220124 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • ATT PM
    • DMS archive index, mongo refresher
    • USB3 extender shark-vis/iLocater iterations
    • safety training
  • This week
    • all-hands presentation
    • shark-nir doc review
    • DMS archive mongo, lit2/webpack
    • PB work preps readiness

YZ

  • Previous:
    • Several MODS issues:
      • mods2data data cleaning (use rm507-1 for help).
      • MODS GUI issues (discussions with Olga and Rick, may have to rewrite certain MODS programs).
      • MODS WCS archive behavior discussion.
      • IT7400.
    • Updated lib-telemetry for a bug fix (pull request merged), next is to find sometime to install it (planning on manually compile and copy the .so file to tcs1).
    • TMS meeting, software and wiki update.
  • Next:
    • Continue some of the MODS work and discussion.
    • A long list of TMS work to be done.
    • Possible TMS test to be scheduled with Heejoo.
    • lib-telemetry update?

DE

  • Vault migration work
    • As far as I can tell we are ready for cutover
    • Jim is not convinced, meeting this morning:
    • Ketiv on standby to handle issues
  • IssueTrak autoassignment
    • Not sure what's up; I would have thought that subtypes would inherit the assignment from the parent.
    • Is it worth opening a ticket with IT?
  • Started work on new Synology to put the Vault backups there
  • Haven't heard back about the Veeam licensing, will ping CDW

GE

  • Last week
    • Fixed numerous minor bugs in the RTD discovered during performance testing.
    • Added a feature to dynamically determine the offset of the bad amplifier on LUCI2.
    • Investigated a deep bug in the startup of the LUCI servers (the check thread is started before registering the server, which can result in premature shutdown), but have not decided how to fix it.
    • Discovered that ssh -YC is better than X2Go for running LUCI simulator on a VM on my local machine.
  • This week
    • Merge the new RTD, and deploy to the observer computers.
    • Performance testing of the new RTD; see the notes in the pull request for details.
    • Begin work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci).
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

PSG

  • Last week
    • Finished coding reset
    • Testing reset
    • Tests were inconclusive, telescope was in wrong configuration.
  • This week
    • Reschedule test.
    • Finsh addressing GCS issue.

BM

  • Last week:
    • LN fastlink test.
    • ARGOS test.
    • OCAM test support dxlbti/dxsoul.
    • max_iteration fix for JCG.
    • merged logging fixes
    • looked into adsec failures from previous week
  • This week:
    • Monday: SOUL SW update tests with Juan Carlos
    • Monday/Tuesday: LBTI SW update tests.
    • Wednesday DEC: LBTI SW checkout (should be tested already)
    • Things to do:
      • Anything that comes up from a or b.
      • SafeSkip ICE communication error.
      • SafeSkip interrupted offsets error.
      • Get AO conf/calib into git.
      • WfsArbitrator timeouts on load.
      • Other GH issues.

SH

  • VM2 PERC card
  • Synology move
    • built RPMS for archive and repository
    • continued to copy data
  • new core switches
  • Hosting and domain for scienceops (DNS changes as requested)
  • Archive stuff
  • iLocater fiber meeting
  • AO switch issue
    • replacements ordered

next:

  • AO switch issue
    • check what is in my office
  • VM2 PERC card
  • new core switches
  • keep moving stuff to Synology
Topic revision: r10 - 31 Jan 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback